Use the store's annual mean and standard deviation for your analysis. Determine the probability that on any1 given day the main street store would have sales

a) greater than 50 atlas tires

b) fewer than 30 atlas tires

c) between 40 atlas tires and 80 atlas tires

Month Main Street Location
20xx Atlas Tires Continental
January 52 12
February 38 22
March 27 26
April 14 0
May 26 0
June 14 8
July 8 16
August 14 4
September 32 12
October 0 24
November 12 0
December 8 18
Main Street Stats
Atlas Tires Continental
Mean Median Mode Deviation Squared Deviation Quartiles Mean Median Mode Deviation Squared Deviation Quartiles
20.4166667 14 14 1st 11 11.83333333 12 0 1st 3
2nd 14 2nd 12
0 0 -20.4167 416.8403 3rd 28.25 0 0 -11.8333 140.0278 3rd 19
8 8 -12.4167 154.1736 4th 52 0 0 -11.8333 140.0278 4th 26
8 8 -12.4167 154.1736 0 0 -11.8333 140.0278
12 12 -8.41667 70.84028 4 4 -7.83333 61.36111
14 14 -6.41667 41.17361 8 8 -3.83333 14.69444
14 14 -6.41667 41.17361 12 12 0.166667 0.027778
14 14 -6.41667 41.17361 12 12 0.166667 0.027778
26 26 5.583333 31.17361 16 16 4.166667 17.36111
27 27 6.583333 43.34028 18 18 6.166667 38.02778
32 32 11.58333 134.1736 22 22 10.16667 103.3611
38 38 17.58333 309.1736 24 24 12.16667 148.0278
52 52 31.58333 997.5069 26 26 14.16667 200.6944
Total 245 2434.917 Total 142 1003.667
Annual Range Standard Deviation Coefficient of Variation Annual Range Standard Deviation Coefficient of Variation
0-52 14.24464 69.77% 0-26 9.14543 77.29%
Answer #1

a) greater than 50 atlas tires

z = (x - µ)/σ = (50 - 20.42)/14.24464 = 2.08

P (z > 50) = 0.0189

b) fewer than 30 atlas tires

z = (x - µ)/σ = (30 - 20.42)/14.24464 = 0.67

P (z < 30) = 0.7494

c) between 40 atlas tires and 80 atlas tires

z = (x - µ)/σ = (40 - 20.42)/14.24464 = 1.37

P (z = 40) = 0.9154

z = (x - µ)/σ = (80 - 20.42)/14.24464 = 4.18

P (z = 80) = 1.0000

Required probability = 1.0000 - 0.9154 = 0.0845